# A2Seek: Towards Reasoning-Centric Benchmark for Aerial Anomaly Understanding

### v1

@@ A.2 Annotation Workflow and Quality Assurance
The A2Seek dataset employs a rigorous multi-level annotation framework to ensure high-quality and comprehensive labeling. The annotation process begins with manually labeling anomaly categories and identifying relevant regions in keyframes, which are then extended to adjacent video frames using Grounded-SAM-2 ravi2024sam2segmentimages ; ren2024grounding . These annotations undergo cross-verification by human annotators to ensure consistency and accuracy. Based on this initial information, prompts were designed to guide a vision-language model in generating detailed annotations, including region-level, frame-level, and segment-level information. The model-generated annotations were further reviewed and refined by professional annotators to ensure alignment with human-labeled categories. To enhance the reliability of the annotations, a multi-stage process was designed, encompassing trigger, diagnose, reasoning, reflection, and seeking phases. Specific formatting rules were employed to align the model outputs with human annotations, ensuring consistency and interpretability. During this process, the model was guided to generate multiple responses, from which professional annotators selected the most appropriate ones for further refinement. The finalized annotations were integrated into a reasoning-centric framework, providing structured explanations that include spatial localization, fine-grained category labels, and causal reasoning paths. This process ensures that the dataset not only supports precise anomaly detection but also facilitates in-depth semantic reasoning. The annotation framework includes precise timestamps for each anomalous event, identifying the specific frame sequences where anomalies occur. Spatial localization is achieved through bounding boxes that accurately mark the positions of anomalous objects, ensuring high precision and reliability in anomaly detection. Additionally, natural language explanations describe the causes and contextual background of anomalies, forming dynamic reasoning paths based on visual scene content. To maintain data quality, low-quality videos were removed, and privacy-sensitive information, such as faces and license plates, was processed to ensure compliance with ethical standards. This meticulous annotation process ensures that the A2Seek dataset provides a robust foundation for developing and evaluating aerial anomaly detection models in complex scenarios.
@@
Figure 7: Four–stage annotation workflow. Step 1 (blue) Temporal annotators use our in-house GUI to mark the start/end frames and class of every anomalous episode, exporting a JSON timeline. Step 2 (salmon) For the first frame of each event, experts draw a bounding box around the anomalous region and supply a natural-language description, yielding single-frame spatial seeds. Step 3 (green) A pretrained tracker propagates each seed through the clip to form full-length trajectories; an automated checker screens the results, triggers relabelling when necessary, and funnels all approved tracks into the spatial-label repository. Step 4 (violet) Vision-language models (VLMs) ingest the temporal tags, spatial tracks, and human captions; via chain-of-thought reasoning, they merge these cues into consolidated frame-level annotations, producing the final per-video spatial label set.
Figure 8: Our dataset contain multiple scene, including those usual scene and unusual scene.
Fifteen domain specialists spent roughly one month creating the A2Seek labels. The pipeline concentrates on three elements—temporal boundaries, spatial trajectories, and textual descriptions—implemented with lightweight in-house tools and a round-robin verification scheme. All outputs are released as COCO-VID-style JSON files.
Temporal Boundaries
We built a browser-based tool that lets an annotator scrub through each clip and press Start/End buttons while selecting one of the 20 anomaly classes. Every video is labelled independently by two experts; a third reviewer resolves mismatches. The final time-stamps are exported to temporal_labels.json together with the flight metadata recorded on location.
Spatial Trajectories
For the first frame of each temporally marked event, an expert draws a bounding box and writes a short noun phrase (e.g., “bike”). Two complementary tracking modes are then launched:
@@
Textual Descriptions
Grounding-SAM’s <Caption> head produces frame-level and clip-level captions. In addition, annotators provide a multi-steps description for every anomaly to capture intent and context. Captions are lightly proof-read before being written to text_labels.json.
@@ A.3 Dataset Features
The anomalous events in the A2Seek dataset can be categorized into long-term, short-term, and instantaneous anomalies based on their duration, as shown in Figure 9(a). Long-term anomalies are generally identifiable from a single frame, although contextual information may sometimes be needed for accurate judgment, and they are often somewhat predictable. Short-term anomalies occur over a brief duration and usually require prior knowledge for detection; some can also be detected from a single frame. Instantaneous anomalies occur very quickly, within just a few frames, making detection challenging. For instance, in cases like theft, the entire act may take place in an instant, which complicates its identification.
(a) Frame-level vs. video-level counts for normal and abnormal segments. The blue bars show total frame counts (left axis), and the orange bars show video counts (right axis).
(b) Distribution of drone-recording attributes. Bars depict number of videos by height category (Low/High), velocity category (Low/Medium/High), and time of day (Morning/Afternoon/Night).
(c) Average frame length of each anomaly type.
Figure 9: Dataset statistics and anomaly duration analysis. (a) Comparison of total frame count versus video count between normal and abnormal segments, illustrating the imbalance in segment lengths. (b) Static attributes of drone recordings: height (H0/H1), velocity (M0/M1/M2), and time-of-day (Morning/Afternoon/Night), showing the coverage of environmental conditions. (c) Average duration (in frames) for each anomaly class, grouped by long-term, short-term, and instantaneous events, highlighting the temporal scale of different anomaly categories.
@@
Scene Related Anomaly
Scene-related anomalies are a significant feature of the A2Seek dataset, encompassing behavior anomalies that depend on specific scene contexts and occlusion anomalies caused by obstructions within the environment. The scene-related anomalies are categorized into two types: cross-video-segment scene-dependent anomalies, where a behavior considered normal in one segment may appear anomalous in another, and scene-dependent anomalies unique to drone perspectives, which are influenced by dynamic changes in viewpoint within the same segment. Occlusion anomalies also arise frequently in drone perspectives due to obstructions such as trees or buildings, posing a unique challenge that requires the model to infer anomalous events by leveraging temporal information from preceding and succeeding frames. These complexities make drone-view anomaly detection significantly different from traditional surveillance settings, emphasizing the importance of effective context-based analysis. These scene-related anomalies are shown in Figure 11 and Figure 12.
@@ Appendix B Reward Design for Aerial Anomaly Understanding
Localization Reward.
The localization reward $R_{\text{loc}}$ is defined as the Intersection over Union (IoU) between the predicted bounding boxes $\hat{B}$ and the ground truth boxes $B^{*}$. This reward encourages the model to focus on the most relevant regions of the input, thereby improving its ability to identify and understand anomalies in complex aerial scenes.
Seeking Reward.
The seeking reward $R_{\text{seek}}$ is designed to incentivize the model to determine the necessity of additional high-resolution context and, when required, to precisely localize the corresponding regions. Formally, it is expressed as $R_{\text{seek}}=\mathbb{I}[\hat{s}=s^{*}]\cdot\text{IoU}(\hat{B},B^{*})$, where $s^{*}$ and $\hat{s}$ are the ground truth and predicted seeking decisions, and $\text{IoU}(\hat{B},B^{*})$ measures the overlap between the predicted bounding box $\hat{B}$ and the ground truth $B^{*}$. This formula rewards the model for directly outputting predictions in simple scenarios, while encouraging it to simulate the process of focusing on suspicious areas by cropping and analyzing these regions for further inspection in complex scenarios.
Length Reward.
[…]
@@ C.1 Self-Correction via Reflection with Reasoning Annotations
$P_{\theta}(\tilde{y}\mid x,y_{0})\propto\exp\left(s_{\theta}(x,\tilde{y})+\lambda\Delta r(x,y_{0},\tilde{y})\right),$ (6)
@@
$\Delta\theta=-\eta\nabla_{\theta}\mathbb{E}[\Delta r(x,y_{0},\tilde{y})].$ (9)
@@
The seeking reward integrates seamlessly with the A-GRPO algorithm, influencing both the reflection reward $\Delta r(x,y_{0},\tilde{y})$ and the policy update. Specifically, the total reward $R(x,y)$ now includes $R_{\text{seek}}$ as a component:
@@ Appendix E Limitations, Open Challenges and Future Directions
We introduce A2Seek—a large-scale, multimodal UAV-view dataset enriched with dense bounding-box tracks and sentence-level descriptions—and A2Seek-R1, a multi-stage RL-fine-tuning framework that not only detects anomalies but also explains where and why they occur. To date, we measure performance with frame-level average precision and mean IoU, but our richly annotated data enable much more: future work will add region-aware metrics (e.g., temporal IoU of predicted vs. ground-truth tubes) and spatio-temporal localization scores that jointly reward correct timing and placement of anomaly predictions. Moreover, the moving UAV perspective—where objects shrink with altitude, perspective shifts continuously, and the camera itself pans—calls for bespoke evaluation criteria that penalize false alarms on distant clutter more heavily and incentivize early anticipation of emerging anomalies.

### v2

@@ A.2 Annotation Workflow and Quality Assurance
Dataset Annotation. The A2Seek dataset employs a rigorous multi-level annotation framework to ensure high-quality and comprehensive labeling. As illustrated in Figure 7, the annotation process begins with manually labeling anomaly categories and identifying relevant regions in keyframes, which are then extended to adjacent video frames using Grounded-SAM-2 ravi2024sam2segmentimages; ren2024grounding. These annotations undergo cross-verification by human annotators to ensure consistency and accuracy. Based on this initial information, prompts were designed to guide a vision-language model in generating detailed annotations, including region-level, frame-level, and segment-level information. The model-generated annotations were further reviewed and refined by professional annotators to ensure alignment with human-labeled categories.
To enhance the reliability of the annotations, a multi-stage process was designed, encompassing trigger, diagnose, reasoning, reflection, and seeking phases. Specific formatting rules were employed to align the model outputs with human annotations, ensuring consistency and interpretability. During this process, the model was guided to generate multiple responses, from which professional annotators selected the most appropriate ones for further refinement. The finalized annotations were integrated into a reasoning-centric framework, providing structured explanations that include spatial localization, fine-grained category labels, and causal reasoning paths. This process ensures that the dataset not only supports precise anomaly detection but also facilitates in-depth semantic reasoning. The annotation framework includes precise timestamps for each anomalous event, identifying the specific frame sequences where anomalies occur. Spatial localization is achieved through bounding boxes that accurately mark the positions of anomalous objects, ensuring high precision and reliability in anomaly detection. Additionally, natural language explanations describe the causes and contextual background of anomalies, forming dynamic reasoning paths based on visual scene content. To maintain data quality, low-quality videos were removed, and privacy-sensitive information, such as faces and license plates, was processed to ensure compliance with ethical standards. This meticulous annotation process ensures that the A2Seek dataset provides a robust foundation for developing and evaluating aerial anomaly detection models in complex scenarios. Fifteen domain specialists spent roughly one month creating the A2Seek labels. The pipeline concentrates on three elements—temporal boundaries, spatial trajectories, and textual descriptions—implemented with lightweight in-house tools (as shown in Figure 8) and a round-robin verification scheme. All outputs are released as COCO-VID-style JSON files.
@@
Temporal Boundaries. We built a browser-based tool that lets an annotator scrub through each clip and press Start/End buttons while selecting one of the 20 anomaly classes. Every video is labelled independently by two experts; a third reviewer resolves mismatches. The final time-stamps are exported to temporal_labels.json together with the flight metadata recorded on location.
Spatial Trajectories. For the first frame of each temporally marked event, an expert draws a bounding box and writes a short noun phrase (e.g., “Bicycle”). Two complementary tracking modes are then launched:
@@
Textual Descriptions. Grounding-SAM’s <Caption> head produces frame-level and clip-level captions. In addition, annotators provide a multi-steps description for every anomaly to capture intent and context. Captions are lightly proof-read before being written to text_labels.json.
Unlike CoT-RAG li2025cot, which models factual relationships between knowledge units, GoT focuses on structuring the reasoning trajectory. This design allows adaptive control of reasoning depth in aerial-view anomaly understanding, where scene complexity and anomaly subtlety vary widely.
Graph-of-Thought (GoT) explicitly structures reasoning through stage-specific tags inserted into the generated sequence: <|Trigger|>, <|Diagnose|>, <|Reasoning|>, <|Reflection|>, <|Seeking|>.
Each stage represents a distinct cognitive operation:
  • Trigger: Scene perception and anomaly signal activation.
  • Diagnose: Hypothesis formation for possible anomaly categories.
  • Reasoning: Contextual analysis and evidence-driven judgment.
  • Reflection: Self-evaluation or consistency check of reasoning.
  • Seeking: Optional localization or reference search for key regions.
Each stage is skippable, allowing GoT to shorten reasoning chains for simple scenes and extend them for complex ones. We provide below two representative GoT annotations illustrating both simple and complex reasoning cases.
GoT Annotation Example (a): Simple Scene (Normal) {
"trigger": "No individuals or objects are present in the scene. The environment consists of a running track, a soccer field, and an adjacent area with no visible activity or anomalies.",
"diagnose": ["Normal"],
"reasoning": "<no need>",
"reflection": "<no need>",
"seeking": "<no need>"
GoT Annotation Example (b): Complex Scene (Lost Item) {
"trigger": "In the sequence of images, there is a noticeable change in the lower left area of the frame. Specifically, a person appears to be walking along a pathway near the edge of the lawn. This individual’s presence was not evident in the initial frames but becomes apparent as the sequence progresses.",
"diagnose": "Possible anomaly categories include ‘Trespassing on Lawn’ or ‘Lost Item’. The appearance of the person near the edge of the lawn could indicate an unusual presence, especially if they are not supposed to be in that area. Alternatively, the person might have dropped something, which would classify as a ‘Lost Item’.",
"reasoning": "The person is seen moving along a designated pathway, which suggests normal pedestrian behavior. However, their proximity to the lawn edge raises suspicion. Upon closer inspection, there is no direct indication of the person interacting with the lawn itself, such as stepping onto it. Instead, the focus shifts to the possibility of a lost item, as the person’s movement does not suggest any abnormal interaction with the environment beyond walking.",
"reflection": "Given the lack of direct interaction with the lawn and the person’s adherence to the pathway, the most plausible explanation is that the person has dropped an item. This aligns with the category ’Lost Item’, as the person’s movement pattern does not suggest trespassing but rather a search for something they may have misplaced.",
"seeking": "[270, 170, 582, 447]"
These examples illustrate how GoT represents reasoning as a semantic graph of thought stages, bridging perception and interpretation in an interpretable and controllable manner.
@@ A.3 Dataset Features
The anomalous events in the A2Seek dataset can be divided into three categories according to their duration: long-term, short-term, and instantaneous anomalies. Long-term anomalies are generally difficult to identify from a single frame and require extended contextual information for accurate judgment; such anomalies are often somewhat predictable, for example, prolonged loitering in restricted areas. Short-term anomalies last for a shorter period and typically require certain prior knowledge for detection, though some can still be recognized from individual frames; for instance, running across a pedestrian walkway may constitute a short-term anomaly. Instantaneous anomalies occur extremely rapidly, usually within only a few frames, making them difficult to detect. For instance, in theft scenarios, the entire action may occur in an instant, thereby complicating the identification process.
@@
Scene-dependent Anomaly
Scene-dependent anomalies are a significant feature of the A2Seek dataset, encompassing behavior anomalies that depend on specific scene contexts and occlusion anomalies caused by environmental obstructions. Scene-dependent anomalies can be categorized into two types: cross-video-segment scene-dependent anomalies, where a behavior that appears normal in one segment may be anomalous in another (as shown in Figure 10), and drone-specific scene-dependent anomalies, influenced by dynamic viewpoint shifts within the same segment (as shown in Figure 11). Occlusion anomalies also frequently occur in drone views due to obstructions such as trees or buildings, posing unique challenges that require models to infer anomalies by leveraging temporal cues from preceding and succeeding frames. These complexities make drone-view anomaly detection substantially different from traditional surveillance settings, highlighting the importance of context-aware analysis.
@@ Appendix B Reward Design for Aerial Anomaly Understanding
Localization Reward. The localization reward $R_{\text{loc}}$ is defined as the Intersection over Union (IoU) between the predicted bounding boxes $\hat{B}$ and the ground truth boxes $B^{*}$. This reward encourages the model to focus on the most relevant regions of the input, thereby improving its ability to identify and understand anomalies in complex aerial scenes.
Seeking Reward. The seeking reward $R_{\text{seek}}$ is designed to incentivize the model to determine the necessity of additional high-resolution context and, when required, to precisely localize the corresponding regions. Formally, it is expressed as $R_{\text{seek}}=\mathbb{I}[\hat{s}=s^{*}]\cdot\text{IoU}(\hat{B},B^{*})$, where $s^{*}$ and $\hat{s}$ are the ground truth and predicted seeking decisions, and $\text{IoU}(\hat{B},B^{*})$ measures the overlap between the predicted bounding box $\hat{B}$ and the ground truth $B^{*}$. This formula rewards the model for directly outputting predictions in simple scenarios, while encouraging it to simulate the process of focusing on suspicious areas by cropping and analyzing these regions for further inspection in complex scenarios.
Length Reward. To address the overthinking phenomenon sui2025stop in simple scenarios, we couple answer correctness with reasoning length $L$ (tokens in the <think> segments). If the answer is correct we favour brevity, setting $R_{d}=\frac{1}{\log(1+L)}$; if it is wrong we encourage elaboration with $R_{d}=\min(L/L_{\max},\,1)$, where $L_{\max}$ equals the model’s maximum output length. The reward is zero whenever either <think> or <answer> is missing.
@@ C.1 Self-Correction via Reflection with Reasoning Annotations
[…]
@@
[…]
@@
[…]
@@ Appendix E Limitations, Open Challenges and Future Directions
We introduce A2Seek—a large‐scale, multimodal UAV-view dataset enriched with dense bounding-box tracks and sentence-level descriptions—and A2Seek-R1, a multi-stage RL-fine-tuning framework that not only detects anomalies but also explains where and why they occur. To date, we measure performance with frame-level average precision and mean IoU, but our richly annotated data enable much more: future work will add region-aware metrics (e.g., temporal IoU of predicted vs. ground-truth tubes) and spatio-temporal localization scores that jointly reward correct timing and placement of anomaly predictions. Moreover, the moving UAV perspective—where objects shrink with altitude, perspective shifts continuously, and the camera itself pans—calls for bespoke evaluation criteria that penalize false alarms on distant clutter more heavily and incentivize early anticipation of emerging anomalies.